Suppose that a random variable X follows a distribution with mean 100 and variance 81. We take a random sample of 240 observations and calculate x. What is the probability that we calculate a sample mean larger than 101?
Answer #1

. r = 9 M=100 r = 81 n=240 :p (X=x) = x-u P(x > 101- 15240 = 15.49 101-100 - 91240 table, the P(x) 101)= 1,7211 Normal distri